# The Transfer Performance of Economic Models

### v4

@@ 2.1. Transfer errors
We will be interested in a large class of measures for transfer performance.
@@ 2.2. The analyst’s problem
We assume that $n>r$; that is, the analyst can collect a larger number of samples than were used by the researcher. In the analyst’s statistical model, the samples $S_{1},S_{2},\dots$ are generated in the following way:
Assumption 1 (Statistical Model).
There is a fixed (but unknown) meta-distribution $\mu\in\Delta(\mathcal{P}\times\mathbb{N})$ over joint distributions $\mathcal{P}\equiv\Delta(\mathcal{X}\times\mathcal{Y})$ and sample sizes $\mathbb{N}$, where each sample $S_{d}$ is generated by first drawing a distribution and sample size $(P_{d},m_{d})\sim\mu$, and then independently drawing $m_{d}$ observations $(x,y)$ from $P_{d}$. [^8] All of our results extend unchanged if samples from the different domains are ex-ante exchangeable.
In Examples 1-3, this assumption implies that (from the analyst’s perspective) the subject pools, choice frames, or choice menus differentiating the samples are themselves drawn i.i.d. from a fixed distribution. Assumption 8 is standard in conformal inference (Vovk et al., 2005), permutation testing Romano (1990), and randomization inference (Ritzwoller et al., 2024), and can also be understood as a version of cluster sampling (Liang and Zeger, 1986; Bugni et al., 2023). [^9] If framed in this way, the analyst’s goal is to do predictive inference for new clusters. When $\mu$ assigns probability $1$ to a single distribution in $p\in\mathcal{P}$ or when $\mu$ assigns probability $1$ to $m=1$, this reduces to i.i.d. sampling of observations from a fixed joint distribution, but our focus is on settings where neither of these is the case. In contrast, the literature on external validity (see Section 1.1) instead typically assumes that the distributions governing behavior in different domains are close in some distance metric (Adjaho and Christensen, 2023), share a common support over $\mathcal{X}$ or $\mathcal{Y}$ (Sahoo et al., 2022; Lei et al., 2023), or can be estimated using background covariates (Tipton and Olsen, 2018). Relative to these assumptions, our approach has the advantage of allowing for arbitrary and unknown relationships between the realized distributions governing domains, but it rules out ex-ante predictable patterns in how the joint distribution varies across samples (such as time trends). In Section 3.2, we extend our main results to the case where the test samples have a different distribution than the training samples. Appendix R.1 further extends our results to a model in which even the training samples are non-identically distributed.
The analyst models the researcher’s set $\mathcal{T}$ of $r$ training domains as drawn uniformly over all subsets of $\{1,\dots,n\}$ of size $r$. We use $\mathbb{T}$ to mean the random variable whose realization is $\mathcal{T}$, so that $S_{\mathbb{T}}=(S_{d})_{d\in\mathbb{T}}$ is the researcher’s (random) vector of training samples. The target domain (on which predictions will be made) is a final sample $S_{n+1}$, which, unlike the metadata, is not observed by the analyst. The quantity of interest is $e_{\mathbb{T},n+1}$, i.e., the random transfer error when the researcher extrapolates predictions from $S_{\mathbb{T}}$ to $S_{n+1}$. Figure 1 depicts this transfer error for the model transfer specification of (2.1).
@@
The analyst’s goal is to develop forecast intervals for the transfer error $e_{\mathbb{T},n+1}$, i.e., interval-valued functions of the meta-data $\mathbb{M}$ which cover $e_{\mathbb{T},n+1}$ with the prescribed probability, regardless of the distribution $\mu$ that governs samples across domains.
2.3. Our procedure
The analyst does not observe the target sample $S_{n+1}$, but (thanks to Assumption 8) can employ the observed samples in the metadata as surrogates for the unseen target sample. As before, let $e_{\mathcal{T},d}^{\mathbf{M}}$ denote the (observed) transfer error from any selection of training samples $\mathcal{T}\subseteq\{1,\dots,n\}$ to any surrogate target sample $d\in\{1,\dots,n\}\backslash\mathcal{T}$ from the metadata (where we now make the dependence of this quantity on $\mathbf{M}$ explicit). We use $\mathbb{T}_{r+1,n}$ to denote the set of $\frac{n!}{(n-r-1)!}$ unique pairs $(\mathcal{T},d)$ that can be constructed in this way. Then
@@ 2.3. Our procedure
Our forecast interval for the transfer error on the target sample is $[\underline{e}_{\tau}^{\mathbf{M}},\overline{e}_{\tau}^{\mathbf{M}}]$; we show in the subsequent section that this is a valid forecast interval.
3. Theoretical Results
3.1. Main Results
We first prove that $[\underline{e}_{\tau}^{\mathbf{M}},\overline{e}_{\tau}^{\mathbf{M}}]$ is indeed a valid forecast interval. Since in many applications only a limited number of domains will be observed, we consider only finite-sample results in this paper.
@@ 2 $e_{2,1}$ddots $\vdots$ $\vdots$
3.2. Relaxing the i.i.d. Assumption
Our results so far assume that the distributions governing the different samples $S_{d}$ are themselves independent and identically distributed. This assumption is not always appropriate. For example, suppose variation in domains corresponds to variation over locations, and the samples in the metadata (but not the target sample) are from experiments run at locations chosen by experimenters. If there is selection bias over where experiments were run—for example, if the observed sites were chosen based on characteristics correlated with effect sizes (as Allcott (2015) found in the Opower energy conservation experiments)—then it may be that the the target sample has fundamentally different properties from anything that is observed in the metadata. We thus now relax Assumption 8 to allow the distribution governing the training samples and the distribution governing the target sample to be drawn from different meta-distributions. (Appendix R.1 further extends Corollary 1 to also allow the training samples to be non-identically distributed. We focus on this simpler version in the main text for ease of exposition.)
Specifically, suppose that the analyst’s metadata consists of samples $S_{1},\dots,S_{n}\sim_{iid}\mu$ as in our main model, but $S_{n+1}$ is independently drawn from some other density $\nu$. Let
@@
3.2.1. The analyst knows the likelihood ratio $\omega$
We again construct a forecast interval for $e_{\mathbf{T},n+1}$ using the pooled sample of transfer errors across samples in the metadata, that is, $\left\{e^{\mathbf{M}}_{\mathcal{T},d}:(\mathcal{T},d)\in\mathbb{T}_{r+1,n}\right\}$. Different from the previous section, we no longer assign uniform weights to each $e^{\mathbf{M}}_{T,d}$. Intuitively, under our previous i.i.d. assumption, each sample in the metadata was equally representative of the training and target distributions, but in this relaxed model whether a sample $S_{d}$ is more representative of the training or testing distribution depends on its relative likelihood under $\nu$ and $\mu$.
@@ 3.2.1. The analyst knows the likelihood ratio $\omega$
3.2.2. The analyst does not know $\omega$ but can bound it
We can again extend our results when the analyst does not know the likelihood ratio function $\omega$ precisely, but knows that it admits an upper and lower bound, as in sensitivity analysis (Rosenbaum, 2005).
@@ 3.2.2. The analyst does not know $\omega$ but can bound it
3.2.3. The analyst knows nothing about $\omega$
Finally, we provide two ways for comparing the transferability of two models $i=1,2$ when the analyst cannot bound $\omega$. We do not provide formal results about these orders, but show that they have bite in our subsequent application (see Section 4). Let $\overline{e}_{i,\tau}^{\mathbf{M}}(\Gamma)$ and $\underline{e}_{i,\tau}^{\mathbf{M}}(\Gamma)$ denote the worst case bounds for model $i$.
@@ 4. Application
To illustrate our methods, we evaluate the transferability of predictions of certainty equivalents for binary lotteries, where the domains correspond to different subject pools (among other potential differences). We focus on this application for several reasons: First, since it is one of the oldest problems in microeconomics, there are many public data sources that we can use to construct our metadata. Second, the associated economic models have been extensively examined from the perspective of predictive performance (Harless and Camerer, 1994; Hey and Orme, 1994; Bruhin et al., 2010; Bernheim and Sprenger, 2020), and recent work evaluates how well these models predict relative to black box algorithms (Peysakhovich and Naecker, 2017; Plonsky et al., 2019; Fudenberg et al., 2022). Finally, as Einav et al. (2012) points out, given how models of risk preferences are often used in practice, it is also important to evaluate how well they transfer across domains. We thus view this application as a natural one with which to illustrate our methods.
@@ 4.4. Transfer error
We use the results in Section 3 to construct forecast intervals for the two specifications of transfer error defined in (2.1) and(2.2), which we will subsequently call raw transfer error and transfer shortfall respectively. We also consider another normalization of the raw transfer error with respect to a proxy for the best achievable error on the target sample. Let $m\in\mathcal{M}$ index a set of models that each prescribe rules $f^{m}$ for mapping data to prediction rules. Then transfer shortfall
@@
In our meta-data there are $n=44$ domains, and we choose $r=1$ of these to use as the training domainm which corresponds to the question, “If the researcher draws one domain at random, and then tries to generalize to another domain, how well will they do?” Figure 4 displays two-sided forecast intervals for transfer performance, transfer deterioration, and transfer shortfall (where $R$ includes all decision rules shown in the figure). These forecast intervals use $\tau=0.95$, so the upper bound of the forecast interval is the 95th percentile of the pooled transfer errors (across choices of the training and test domains), and the lower bound of the forecast interval is the 5th percentile of the pooled transfer errors. (See Table 5 in Appendix S.3 for the exact numbers.) Applying Proposition 1, these are 86% forecast intervals. Choosing larger $\tau$ results in wider forecast intervals that have higher coverage levels, and we report some of these alternative forecast intervals in Online Appendix S.4, including a 96% forecast interval.
@@
We next use our theoretical results from Section 3.2 to study the consequences of relaxing the i.i.d.  assumption in our comparison of CPT$(\alpha,\beta,\delta,\gamma)$ and RF. Since the main differences observed above concerned the upper bounds of our forecast intervals, we limit attention to $\tau\geq 0.5,$ and compare the methods in terms of worst-case and everywhere upper-dominance with respect to all three measures of the transfer performance. These results are summarized in Table 2.
@@ 4.5. Do black boxes transfer poorly because they are too flexible?
One tempting explanation of why the black box algorithms transfer less well is that they may overfit to idiosyncratic details of the training samples that do not generalize across subject pools. For example, suppose some subject pools tend to value lotteries depending on the specific digits they contain. [^20] For example, Fortin et al. (2014) find that in neighborhoods with a higher than average percentage of Chinese residents, homes with address numbers ending in “4” are sold at a 2.2% discount and those ending in “8” are sold at a 2.5% premium. This regularity could not be captured by the economic models, because they do not include parameters for individual digits, but could be learned by a random forest algorithm. This would lead the random forest to have better within-domain prediction for those subject pools, but worse transfer performance if this regularity does not generalize across subject pools.
@@ 5. Conclusion
Our measures of transfer error quantify how well a model’s performance on one domain extrapolates to other domains. We applied these measures to show that the predictions of expected utility theory and cumulative prospect theory outperform those of black box models on out-of-domain tests, even though the black boxes generally have lower out-of-sample prediction errors within a given domain. The relatively worse transfer performance of the black boxes seems to be because the black box algorithms have not identified structure that is commonly shared across domains, and thus cannot effectively extrapolate behavior from one set of features to another. Our finding that the economic models transfer better supports the intuition that economic models can recover regularities that are general across a variety of domains.
@@ A.3. Proof of Theorem 1 and Corollary 1
Our main results focus on forecasting realized transfer errors, which is useful when we want to know the range of plausible errors in transferring a given model to a new domain. We now complement those results with procedures for inference focused on population quantities: Section Q.2 provides confidence intervals for quantiles of the transfer error distribution, and Section Q.3 provides a confidence interval for the expected transfer error. Since these quantities can be perfectly recovered given data from an infinite number of domains, we expect the lengths of these intervals to vanish as the number of observed domains grows large, unlike the forecast intervals from Section 3.
@@
Appendix R  Supplementary Material to Section 3.2
@@
We briefly describe the individual samples in our meta-data. There are 44 domains in total.
@@
We now consider an alternative definition of domains, with each of the 14 papers representing a different domain. This changes the content of the i.i.d. assumption imposed in Section 3, where we now assume that samples are i.i.d. across papers, but may be dependent across subject pools within the same paper. We repeat our main analysis and report 78% two-sided forecast intervals in Figure 10. These intervals are qualitatively similar to those reported in Figure 4.
@@
Here we consider an alternative choice for the number of training domains, setting $r=3$ instead of $r=1$. This corresponds to randomly choosing 3 of the 44 domains to be the training domains, finding the best prediction rule for this pooled data, and using the estimated prediction rule to predict the remaining 41 samples. For this analysis we use domain cross-validation to select tuning parameters for the black box algorithms, as described in Example 10.

### v5

@@ 2.1. Transfer errors
Our results apply to the following class of measures for transfer performance.
@@ 2.2. The analyst’s problem
We assume that $n>r$; that is, the analyst can collect a larger number of samples than were used by the researcher. The analyst models the researcher’s set $\mathcal{T}$ of $r$ training domains as drawn uniformly over all subsets of $\{1,\dots,n\}$ of size $r$. We use $\mathbb{T}$ to mean the random variable whose realization is $\mathcal{T}$, so that $S_{\mathbb{T}}=(S_{d})_{d\in\mathbb{T}}$ is the researcher’s (random) vector of training samples. The target domain (on which predictions will be made) is a final sample $S_{n+1}$, which, unlike the metadata, is not observed by the analyst. The quantity of interest is $e_{\mathbb{T},n+1}$, i.e., the random transfer error when the researcher extrapolates predictions from $S_{\mathbb{T}}$ to $S_{n+1}$. Figure 1 depicts this transfer error for the model transfer specification of (2.1).
@@
The analyst’s goal is to develop forecast intervals for the transfer error $e_{\mathbb{T},n+1}$, i.e., interval-valued functions of the meta-data $\mathbb{M}$ which cover $e_{\mathbb{T},n+1}$ with the prescribed probability, regardless of the distribution $\mu$ that governs samples across domains. Section 3 provides these intervals for a baseline setting in which domains are exchangeable, and Section 4 generalizes them to a setting in which the observed domains are systematically different from the target domain.
3. Baseline Setting: Exchangeability
We begin with a baseline setting in which the different domains are governed by exchangeable distributions. Specifically, we suppose that in the analyst’s statistical model, the samples $S_{1},S_{2},\dots$ are generated in the following way:
Assumption 1.
There is a fixed (but unknown) meta-distribution $\mu\in\Delta(\mathcal{P}\times\mathbb{N})$ over joint distributions $\mathcal{P}\equiv\Delta(\mathcal{X}\times\mathcal{Y})$ and sample sizes $\mathbb{N}$, where each sample $S_{d}$ is generated by first drawing a distribution and sample size $(P_{d},m_{d})\sim\mu$, and then independently drawing $m_{d}$ observations $(x,y)$ from $P_{d}$. [^7] All of our results extend unchanged if samples from the different domains are ex-ante exchangeable rather than i.i.d..
In Examples 1-3, this assumption implies that (from the analyst’s perspective) the subject pools, choice frames, or choice menus differentiating the samples are themselves drawn i.i.d. from a fixed distribution. Assumption 7 is standard in conformal inference (Vovk et al., 2005), permutation testing Romano (1990), and randomization inference (Ritzwoller et al., 2024). [^8] It can also be understood as a Bayesian hierarchical model (Meager, 2019, 2022) or a version of cluster sampling (Liang and Zeger, 1986; Bugni et al., 2023).If framed in this way, the analyst’s goal is to do predictive inference for new clusters. When $\mu$ assigns probability $1$ to a single distribution in $p\in\mathcal{P}$ or when $\mu$ assigns probability $1$ to $m=1$, this reduces to i.i.d. sampling of observations from a fixed joint distribution, but our focus is on settings where neither of these is the case. In contrast, the literature on external validity (see Section 1.1) typically assumes that the distributions governing behavior in different domains are close in some distance metric (Adjaho and Christensen, 2022), share a common support over $\mathcal{X}$ or $\mathcal{Y}$ (Sahoo et al., 2022; Lei et al., 2023), or can be estimated using background covariates (Tipton and Olsen, 2018). Relative to these assumptions, our approach has the advantage of allowing for arbitrary and unknown relationships between the realized distributions governing domains, but it rules out ex-ante predictable patterns in how the joint distribution varies across samples (such as time trends).
Section 3.1 presents our forecast intervals for this setting, and Section 3.2 proves the validity and tightness of these intervals.
3.1. Baseline procedure
Thanks to Assumption 7, the observed samples in the metadata, $\{S_{1},\dots,S_{n}\}$, can act as surrogates for the unseen target sample, $S_{n+1}$. As before, let $e_{\mathcal{T},d}^{\mathbf{M}}$ denote the (observed) transfer error from any selection of training samples $\mathcal{T}\subseteq\{1,\dots,n\}$ to any surrogate target sample $d\in\{1,\dots,n\}\backslash\mathcal{T}$ from the metadata (where we now make the dependence of this quantity on $\mathbf{M}$ explicit). We use $\mathbb{T}_{r+1,n}$ to denote the set of $\frac{n!}{(n-r-1)!}$ unique pairs $(\mathcal{T},d)$ that can be constructed in this way. Then
@@ 3.1. Baseline procedure
Our proposed forecast interval for the transfer error on the target sample is $[\underline{e}_{\tau}^{\mathbf{M}},\overline{e}_{\tau}^{\mathbf{M}}]$.
3.2. Results
We first prove that $[\underline{e}_{\tau}^{\mathbf{M}},\overline{e}_{\tau}^{\mathbf{M}}]$ is indeed a valid forecast interval.
@@ 2 $e_{2,1}$ddots $\vdots$ $\vdots$
4. Beyond Exchangeability
Our results so far assume that the distributions governing the different samples $S_{d}$ are themselves independent and identically distributed. This assumption is not always appropriate. For example, suppose variation in domains corresponds to variation over locations, and the samples in the metadata (but not the target sample) are from experiments run at locations chosen by experimenters. If there is selection bias over where experiments are run—for example, if the observed sites are chosen based on characteristics correlated with effect sizes (as Allcott (2015) found in the Opower energy conservation experiments)—then it may be that the the target sample has fundamentally different properties from anything that is observed in the metadata. We thus now relax Assumption 7 to allow the distribution governing the training samples and the distribution governing the target sample to be drawn from different meta-distributions. Specifically, suppose that the analyst’s metadata consists of samples $S_{1},\dots,S_{n}\sim_{iid}\mu$ as in our main model, but $S_{n+1}$ is independently drawn from some other density $\nu$. Let
@@ 4. Beyond Exchangeability
The following subsections consider successively weaker assumptions about what the analyst knows about the likelihood ratio $\omega$. Sections 4.1 and 4.2 respectively generalize our previous results when the analyst either knows $\omega$ or can bound it. Section 4.3 provides two ways of comparing the generalizability of models in environments where the analyst knows nothing about $\omega$.
4.1. The analyst knows the likelihood ratio $\omega$
We again construct a forecast interval for $e_{\mathbf{T},n+1}$ using the pooled sample of transfer errors across samples in the metadata, that is, $\left\{e^{\mathbf{M}}_{\mathcal{T},d}:(\mathcal{T},d)\in\mathbb{T}_{r+1,n}\right\}$. Different from the previous section, we no longer assign uniform weights to each $e^{\mathbf{M}}_{T,d}$. Intuitively, under our previous i.i.d. assumption, each sample in the metadata was equally representative of the training and target distributions, but in this relaxed model, whether a sample $S_{d}$ is more representative of the training or testing distribution depends on its relative likelihood under $\nu$ and $\mu$.
@@ 4.1. The analyst knows the likelihood ratio $\omega$
4.2. The analyst does not know $\omega$ but can bound it
We next extend our results when the analyst does not know the likelihood ratio function $\omega$ precisely, but—as in the literature on sensitivity analysis (Rosenbaum, 2005)—knows that it admits an upper and lower bound.
@@ 4.2. The analyst does not know $\omega$ but can bound it
4.3. The analyst knows nothing about $\omega$
Finally, we provide two ways for comparing the transferability of two models $i=1,2$ when the analyst cannot bound $\omega$. We do not provide formal results about these orders, but show that they have bite in our subsequent application (see Section 5).
Let $\overline{e}_{i,\tau}^{\mathbf{M}}(\Gamma)$ and $\underline{e}_{i,\tau}^{\mathbf{M}}(\Gamma)$ again denote the worst case bounds for model $i$, as defined in (4.2).
@@ 5. Application
To illustrate our methods, we evaluate the transferability of predictions of certainty equivalents for binary lotteries across domains that have different subject pools and also differ in other ways. We focus on this application for several reasons: First, there are many public data sources that we can use to construct our metadata. Second, the associated economic models have been extensively examined from the perspective of predictive performance (Harless and Camerer, 1994; Hey and Orme, 1994; Bruhin et al., 2010; Bernheim and Sprenger, 2020), and recent work evaluates how well these models predict relative to black box algorithms (Peysakhovich and Naecker, 2017; Plonsky et al., 2019; Fudenberg et al., 2022). Finally, as Einav et al. (2012) points out, given how models of risk preferences are often used in practice, it is also important to evaluate how well they transfer across domains. We thus view this application as a natural setting to illustrate our methods.
@@ 5.4. Transfer error
We use the results in Section 3.2 to construct forecast intervals for the two specifications of transfer error defined in (2.1) and(2.2), which we will subsequently call raw transfer error and transfer shortfall respectively. We also consider another normalization of the raw transfer error with respect to a proxy for the best achievable error on the target sample. Let $m\in\mathcal{M}$ index a set of models that each prescribe rules $f^{m}$ for mapping data to prediction rules. Then transfer shortfall
@@
In our meta-data there are $n=44$ domains, and we choose $r=1$ of these to use as the training domain which corresponds to the question, “If the researcher draws one domain at random, and then tries to generalize to another domain, how well will they do?” Figure 4 displays two-sided forecast intervals for transfer performance, transfer deterioration, and transfer shortfall (where $R$ includes all decision rules shown in the figure). These forecast intervals use $\tau=0.95$, so the upper bound of the forecast interval is the 95th percentile of the pooled transfer errors (across choices of the training and test domains), and the lower bound of the forecast interval is the 5th percentile of the pooled transfer errors. (See Table 5 in Appendix S.3 for the exact numbers.) Applying Proposition 1, these are 86% forecast intervals. Choosing larger $\tau$ results in wider forecast intervals that have higher coverage levels, and we report some of these alternative forecast intervals in Online Appendix S.4, including a 96% forecast interval.
@@
We next use our theoretical results from Section 4 to study the consequences of relaxing the i.i.d.  assumption in our comparison of CPT$(\alpha,\beta,\delta,\gamma)$ and RF. Since the main differences observed above concerned the upper bounds of our forecast intervals, we limit attention to $\tau\geq 0.5,$ and compare the methods in terms of worst-case and everywhere upper-dominance with respect to all three measures of transfer performance. These results are summarized in Table 2.
@@ 5.5. Do black boxes transfer poorly because they are too flexible?
One tempting explanation of why the black box algorithms transfer less well is that they may overfit to idiosyncratic details of the training samples that do not generalize across subject pools. For example, suppose some subject pools tend to value lotteries depending on the specific digits they contain. [^19] For example, Fortin et al. (2014) find that in neighborhoods with a higher than average percentage of Chinese residents, homes with address numbers ending in “4” are sold at a 2.2% discount and those ending in “8” are sold at a 2.5% premium. This regularity could not be captured by the economic models, because they do not include parameters for individual digits, but could be learned by a random forest algorithm. If so, the random forest would have better within-domain prediction for those subject pools, but worse transfer performance (assuming this regularity does not generalize across subject pools).
@@ 6. Conclusion
Our measures of transfer error quantify how well a model’s performance on one domain extrapolates to other domains. We applied these measures to show that the predictions of expected utility theory and cumulative prospect theory outperform those of black box models on out-of-domain tests, even though the black boxes generally have lower out-of-sample prediction errors within a given domain. The relatively worse transfer performance of the black boxes seems to be because the black box algorithms have not identified structure that is commonly shared across domains, and thus cannot effectively extrapolate behavior from one set of features to another. Our finding that the economic models transfer better supports the intuition that economic models capture regularities that are general across a variety of domains.
One may view our theoretical results, and the statistical assumptions which justify them, from two perspectives. Taken literally, we consider a researcher who has access to data from a small set of domains, and an analyst who is interested in how well the researcher’s procedure extrapolates. To provide guarantees we restrict the ways domains relate to each other; the simplest and strongest of these is that these is that the domains are exchangeable. We also assume that the analyst has access to data from a larger set of domains than the researcher does. One might wonder why the researcher uses a small set of domains for training, when the larger meta dataset is available to the analyst. In particular, the researcher might prefer use the larger cross-domain dataset for training, leaving no “fresh” domains for performance evaluation. Although one could potentially weaken our data requirements to cover such cases by restricting the researcher’s estimator, this would rule out the black box prediction methods that are a focus of our analysis.
Alternatively, one can view our results as a tool for evaluating and comparing the performance of different methods for cross-domain prediction. The starting point for our analysis of a given method is the matrix collecting that method’s cross-domain prediction errors for a set of observed domains. These matrices are rich objects, and it is not obvious how to compare them, but our theoretical analysis points to a natural set of summary statistics: quantiles of the observed error distribution. Focusing on these quantiles, or equivalently on our forecast intervals, provides a theoretically-motivated way to make comparisons across methods.
@@ A.3. Proof of Theorem 1 and Corollary 1
Our main results focus on forecasting realized transfer errors, which is useful when we want to know the range of plausible errors in transferring a given model to a new domain. We now complement those results with procedures for inference focused on population quantities: Section Q.2 provides confidence intervals for quantiles of the transfer error distribution, and Section Q.3 provides a confidence interval for the expected transfer error. Since these quantities can be perfectly recovered given data from an infinite number of domains, we expect the lengths of these intervals to vanish as the number of observed domains grows large, unlike the forecast intervals from Section 3.2.
@@
Appendix R  Supplementary Material to Section 4
@@
We briefly describe the individual samples in our meta-data.
@@
We now consider an alternative definition of domains, with each of the 14 papers representing a different domain. This changes the content of the i.i.d. assumption imposed in Section 3.2, where we now assume that samples are i.i.d. across papers, but may be dependent across subject pools within the same paper. We repeat our main analysis and report 78% two-sided forecast intervals in Figure 10. These intervals are qualitatively similar to those reported in Figure 4.
@@
Here we consider an alternative choice for the number of training domains, setting $r=3$ instead of $r=1$.
This corresponds to randomly choosing 3 of the 44 domains to be the training domains, finding the best prediction rule for this pooled data, and using the estimated prediction rule to predict the remaining 41 samples. For this analysis we use domain cross-validation to select tuning parameters for the black box algorithms, as described in Example 10.